### v1

@@ Abstract
Monocular Depth and Surface Normals Estimation (MDSNE) is crucial for tasks such as 3D reconstruction, autonomous navigation, and underwater exploration. Current methods rely either on discriminative models, which struggle with transparent or reflective surfaces, or generative models, which, while accurate, are computationally expensive. This paper presents a novel deep learning model for MDSNE, specifically tailored for underwater environments, using a hybrid architecture that integrates Convolutional Neural Networks (CNNs) with Transformers, leveraging the strengths of both approaches. Training effective MDSNE models is often hampered by noisy real-world datasets and the limited generalization of synthetic datasets. To address this, we generate pseudo-labeled real data using multiple pre-trained MDSNE models. To ensure the quality of this data, we propose the Depth Normal Evaluation and Selection Algorithm (DNESA), which evaluates and selects the most reliable pseudo-labeled samples using domain-specific metrics. A lightweight student model is then trained on this curated dataset. Our model reduces parameters by 90% and training costs by 80%, allowing real-time 3D perception on resource-constrained devices. Key contributions include: a novel and efficient MDSNE model, the DNESA algorithm, a domain-specific data pipeline, and a focus on real-time performance and scalability. Designed for real-world underwater applications, our model facilitates low-cost deployments in underwater robots and autonomous vehicles, bridging the gap between research and practical implementation.
@@ 4.4 Quantitative Comparison
To fairly evaluate the performance of our proposed model, PoseidonNet, we performed a quantitative comparison against several state-of-the-art methods of similar size for both depth estimation and surface normals estimation, ensuring a fair comparison without including significantly larger models. All baseline model results were obtained from the published articles of the respective authors to ensure a fair and accurate comparison.
Depth Estimation: We evaluated the depth estimation capabilities of PoseidonNet on three diverse underwater datasets: USOD 10K [11], Atlantis [28] , and Sea-thru [1]. These datasets were chosen to assess the model’s performance across various underwater scenarios and its ability to generalise to unseen data.
USOD 10K: As shown in table 2, PoseidonNet significantly outperformed all other methods. It achieved the highest accuracy scores across all three thresholds ($\delta_{1}$, $\delta_{2}$, and $\delta_{3}$) while simultaneously achieving the lowest error rates (Abs Rel, Sq Rel, RMSE, and log10).
Atlantis: Similar to its performance on USOD 10K, PoseidonNet demonstrated superior performance on the Atlantis dataset, table 3, surpassing other methods in both accuracy and error metrics.
Sea-thru: On the Sea-thru (D3) dataset, PoseidonNet again achieved the best results across all evaluation metrics, further validating its effectiveness in accurate depth estimation for diverse underwater scenes. Table Table 4 provides a detailed breakdown of the results, with PoseidonNet consistently achieving the lowest error rates.
Surface Normals Estimation: Evaluating surface normal estimation in underwater environments presents a unique challenge due to the lack of an established benchmark or dataset specifically designed for this task. To address this limitation, we trained and evaluated PoseidonNet on the iBims-1 dataset [14]. Although iBims-1 was not created for underwater images, it serves as a robust benchmark for assessing surface normal estimation techniques. As presented in Table 5, PoseidonNet achieved competitive results on the iBims-1 benchmark. It performed strongly across key metrics like mean and median errors, as well as accuracy thresholds at various angles.
@@ 4.5 Qualitative Comparison
In Figure 6, we present a comparative analysis between our model (PoseidonNet), Marigold [12], and Metric3D-V2 [25], using open-world images. All models were evaluated using their best publicly available versions from the HuggingFace model hub [9]. The results consistently demonstrate that our method delivers more refined and accurate predictions for both depth and surface normals, especially in challenging scenarios where existing models, such as Marigold and Metric3D-V2, tend to exhibit limitations. Our approach demonstrates a noticeable improvement in sharpness and structure preservation, further illustrating its effectiveness in handling real-world imagery.
@@ 4.7 Computational Efficiency
Beyond accuracy, we also analysed PoseidonNet’s computational efficiency, a crucial factor for real-time applications and deployment on resource-constrained devices. As illustrated in Table 8, PoseidonNet demonstrated a notable balance between accuracy and efficiency. It boasts a significantly lower parameter count and requires less memory compared to more complex models like TransDepth. Furthermore, PoseidonNet achieves a high frame rate (FPS), enabling real-time performance with minimal computational resources.
Table 8: Comparison Results for Complexity and FPS of the proposed PoseidonNet and compared methods. On a Linux platform using a single NVIDIA GeForce RTX 4090 GPU
@@ 5 Discussion
The success of our proposed PoseidonNet model stems from its hybrid architecture, which effectively integrates the strengths of CNNs and Transformers. CNNs excel at extracting local features and details, while Transformers are adept at capturing long-range dependencies and global context. This combination is particularly valuable in underwater environments, where variations in lighting, turbidity, and object textures complicate accurate depth and surface normal estimation.
@@ 6 Conclusions
This research presents a practical solution for the estimation of monocular depth and surface normals in underwater environments. By combining a hybrid CNN-Transformer architecture with an innovative pseudo-labelling technique, the proposed model achieves high accuracy while significantly reducing computational demands and reliance on large labelled datasets. PoseidonNet, our hybrid model, is designed for efficient depth and surface normal estimation, with DNESA improving the quality of pseudo-labelled data and enhancing generalisation. This approach bridges the gap between research and real-world application, enabling real-time 3D perception on resource-constrained devices. Our model is particularly suited for underwater tasks such as seafloor mapping, coral reef monitoring, and species population studies. Future work will expand the model’s capabilities for broader marine applications and improve its integration into real-world systems.

### v2

@@ Abstract
Monocular Depth and Surface Normals Estimation (MDSNE) is crucial for tasks such as three-dimensional (3D) reconstruction, autonomous navigation, and underwater exploration. Current methods rely either on discriminative models, which struggle with transparent or reflective surfaces, or generative models, which, while accurate, are computationally expensive. This paper presents a novel deep learning model for Monocular Depth and Surface Normals Estimation (MDSNE), specifically tailored for underwater environments, using a hybrid architecture that integrates Convolutional Neural Networks (CNNs) with Transformers, leveraging the strengths of both approaches. Training effective MDSNE models is often hampered by noisy real-world datasets and the limited generalization of synthetic datasets. To address this, we generate pseudo-labeled real data using multiple pre-trained MDSNE models. To ensure the quality of this data, we propose the Depth Normal Evaluation and Selection Algorithm (DNESA), which evaluates and selects the most reliable pseudo-labeled samples using domain-specific metrics. A lightweight student model is then trained on this curated dataset. Our model reduces parameters by 90% and training costs by 80%, allowing real-time three-dimensional (3D) perception on resource-constrained devices. Key contributions include: a novel and efficient MDSNE model, the DNESA algorithm, a domain-specific data pipeline, and a focus on real-time performance and scalability. Designed for real-world underwater applications, our model facilitates low-cost deployments in underwater robots and autonomous vehicles, bridging the gap between research and practical implementation.
@@ 4.4 Quantitative Comparison
To fairly evaluate the performance of our proposed model, AquaMDS, we performed a quantitative comparison against several state-of-the-art methods of similar size for both depth estimation and surface normals estimation, ensuring a fair comparison without including significantly larger models. All baseline model results were obtained from the published articles of the respective authors to ensure a fair and accurate comparison.
Depth Estimation: We evaluated the depth estimation capabilities of AquaMDS on three diverse underwater datasets: USOD 10K [11], Atlantis [28] , and Sea-thru [1]. These datasets were chosen to assess the model’s performance across various underwater scenarios and its ability to generalise to unseen data.
USOD 10K: As shown in table 2, AquaMDS significantly outperformed all other methods. It achieved the highest accuracy scores across all three thresholds ($\delta_{1}$, $\delta_{2}$, and $\delta_{3}$) while simultaneously achieving the lowest error rates (Abs Rel, Sq Rel, RMSE, and log10).
Atlantis: Similar to its performance on USOD 10K, AquaMDS demonstrated superior performance on the Atlantis dataset, table 3, surpassing other methods in both accuracy and error metrics.
Sea-thru: On the Sea-thru (D3) dataset, AquaMDS again achieved the best results across all evaluation metrics, further validating its effectiveness in accurate depth estimation for diverse underwater scenes. Table Table 4 provides a detailed breakdown of the results, with AquaMDS consistently achieving the lowest error rates.
Surface Normals Estimation: Evaluating surface normal estimation in underwater environments presents a unique challenge due to the lack of an established benchmark or dataset specifically designed for this task. To address this limitation, we trained and evaluated AquaMDS on the iBims-1 dataset [14]. Although iBims-1 was not created for underwater images, it serves as a robust benchmark for assessing surface normal estimation techniques. As presented in Table 5, AquaMDS achieved competitive results on the iBims-1 benchmark. It performed strongly across key metrics like mean and median errors, as well as accuracy thresholds at various angles.
@@ 4.5 Qualitative Comparison
In Figure 6, we present a comparative analysis between our model (AquaMDS), Marigold [12], and Metric3D-V2 [25], using open-world images. All models were evaluated using their best publicly available versions from the HuggingFace model hub [9]. The results consistently demonstrate that our method delivers more refined and accurate predictions for both depth and surface normals, especially in challenging scenarios where existing models, such as Marigold and Metric3D-V2, tend to exhibit limitations. Our approach demonstrates a noticeable improvement in sharpness and structure preservation, further illustrating its effectiveness in handling real-world imagery.
@@ 4.7 Computational Efficiency
Beyond accuracy, we also analysed AquaMDS’s computational efficiency, a crucial factor for real-time applications and deployment on resource-constrained devices. As illustrated in Table 8, AquaMDS demonstrated a notable balance between accuracy and efficiency. It boasts a significantly lower parameter count and requires less memory compared to more complex models like TransDepth. Furthermore, AquaMDS achieves a high frame rate (FPS), enabling real-time performance with minimal computational resources.
Table 8: Comparison Results for Complexity and FPS of the proposed AquaMDS and compared methods. On a Linux platform using a single NVIDIA GeForce RTX 4090 GPU
@@ 5 Discussion
The success of our proposed AquaMDS model stems from its hybrid architecture, which effectively integrates the strengths of CNNs and Transformers. CNNs excel at extracting local features and details, while Transformers are adept at capturing long-range dependencies and global context. This combination is particularly valuable in underwater environments, where variations in lighting, turbidity, and object textures complicate accurate depth and surface normal estimation.
@@ 6 Conclusions
This research presents a practical solution for the estimation of monocular depth and surface normals in underwater environments. By combining a hybrid CNN-Transformer architecture with an innovative pseudo-labelling technique, the proposed model achieves high accuracy while significantly reducing computational demands and reliance on large labelled datasets. AquaMDS, our hybrid model, is designed for efficient depth and surface normal estimation, with DNESA improving the quality of pseudo-labelled data and enhancing generalisation. This approach bridges the gap between research and real-world application, enabling real-time 3D perception on resource-constrained devices. Our model is particularly suited for underwater tasks such as seafloor mapping, coral reef monitoring, and species population studies. Future work will expand the model’s capabilities for broader marine applications and improve its integration into real-world systems.